OpenStudy (anonymous):

Solve by using the quadratic formula. x2 − 6x + 39 = 0

OpenStudy (anonymous):

a = 1 b = -6 c = 39 so now the formula for quadratics is \[-b + \sqrt{b^2 - 4ac} / 2a\]

OpenStudy (anonymous):

so now substitute and valuate

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

someone help me please

jhonyy9 (jhonyy9):

x_1,_2=(6+/- sqrt(36-156))/2 =(6+/- sqrt(-120))/2 =(6+/- 2isqrt(30))/2 = =3+/- isqrt30